Clear Shine window cleaner is 12% alcohol and Sunstream window cleaner is 30% alcohol. How much of each should be used to make 90 oz of a cleaner that is 20% alcohol.


Answer by jorel1380(3719)

90x.2=18
m+n=90
.12m+.3n=18
12m+30n=1800
12m+12n=1080
18n=720
n=40
m=50
You need 50 oz. of Clear Shine and 40 oz. of Sunstream to make your mixture..